You are playing tetherball with a friend and hit the ball so that it begins to travel in a circular horizontal path. If the ball is 1.2 meters from the pole, has a speed of 3.7 m/s, a mass of 0.42 kilograms, and its (weightless) rope makes a 49° angle with the pole, find the tension force that the rope exerts on the ball just after you hit it.
1
Expert's answer
2019-08-19T09:51:34-0400

Consider the schematic figure:



The 2nd Newton's law applied to the current problem states:

"m\vec{a} = m \vec{g} + \vec{T}"

According to the conditions given, the centripetal force is horizontal. Thus, making the horizontal projection of the equation above, we obtain:

"m a_{c.p.} = T \sin{\alpha} \quad \Rightarrow \quad T = \frac{m a_{c.p.}}{\sin{\alpha}}"

The centripetal acceleration can be calculated as:

"a_{c.p.} = \frac{v^2}{r} = \frac{v^2}{l \sin{\alpha}}"

Finally, we derive:

"T = \frac{m v^2}{l \sin^2{\alpha}}"

Substituting the numerical values, we obtain:

"T = \frac{0.42 \cdot 3.7^2}{1.2 \cdot \sin^2{49^\circ}} \approx 8.4 \, N"


Answer: 8.4 N